OpenStudy (anonymous):

Solve this application using logarithms. At what interest rate (to the nearest hundredth of a percent) compounded annually will money in savings double in five years?

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (elonasushchik):

2 = 1(1 + r/100)^5 2 = (1 + r/100)^5 2^(1/5) = 1 + r/100) 0.2log(10)2 = (100 + r) / 100 = log(10)(100 + r ) - log(10)100 0.060206 = log(10)(!00 + r) - 2 2.060206 = log(10)(100 + r) Antilog . 114.8698 = 100 + r 14.6898 = r r = 14.87% <----answer

OpenStudy (anonymous):

sorry wrong question

OpenStudy (anonymous):

Solve this application using logarithms. Refer to the examples above. How much principal, to the nearest dollar, should you invest at 4% in order to have $3,000 for a vacation to Europe in 3 years?

OpenStudy (elonasushchik):

3000 = i(o)(1 + 4/100)^3 I(o) = 3000 / 1.04^3 i(o) = $2666.99 <----ANSWER Medal and fan?! Thx

OpenStudy (anonymous):

Solve this application using logarithms. How long will money in savings take to double at 5% interest compounded annually? Round the answer to the nearest hundredth of a year.

OpenStudy (elonasushchik):

2 = 1(1+ 5/100)^n 2 = (1.05)^n log(10)2 = nlog(10)1.05 0.301030 = n0.021189 n = 0.30103/0.021189 = 14.20689 yrs. n = 14.21 yrs. <------Answer Medal thx

OpenStudy (anonymous):

Solve this application using logarithms. Refer to the examples above. At his son's birth, a man invested $2,000 in savings at 6% for his son's college education. Approximately how much, to the nearest dollar, will be available in 19 years? ≈ $a0 Because of inflation, money is depreciating in real value. If the rate of inflation is 8% a year, what is the current real value, to the nearest dollar, from the preceding problem above? After 19 years, did the father and son have more or less money, as measured by current real value, than they did at the son's birth? real value ≈ $a1 Type X in front of the correct statement. a2They had more money. a3They had less money.

OpenStudy (elonasushchik):

The first one

OpenStudy (anonymous):

all of them

OpenStudy (anonymous):

what about the other questions?

OpenStudy (elonasushchik):

@ganeshie8 please help him.. Thx

ganeshie8 (ganeshie8):

@thomasfire please limit to one question per post

OpenStudy (anonymous):

At his son's birth, a man invested $2,000 in savings at 6% for his son's college education. Approximately how much, to the nearest dollar, will be available in 19 years?

ganeshie8 (ganeshie8):

close this question to open a new one.. that way your new question gets bumped to top and others see :)